OpenStudy (anonymous):

sqrt (1/12) - sqrt (1/27)

OpenStudy (anonymous):

When you have the sqrt of a fraction, it is the sqrt of the numerator divided by the sqrt of the denominator. Can you convert both expressions to a fraction with denominator 6?

OpenStudy (anonymous):

Yes I can. The first thing you want to do is get the 1's out of the squareroot. \[\sqrt{a/b}=\sqrt{a}/\sqrt{b}\] So \[\sqrt{1/b}=\sqrt{1}/\sqrt{b}=1/\sqrt{b}\]

OpenStudy (anonymous):

o okay

OpenStudy (anonymous):

Does that make sense or did I use too many symbols and make it too long winded?

OpenStudy (anonymous):

no it makes sense thank u!

OpenStudy (anonymous):

The next thing to do is to simplify \(\sqrt{12},\sqrt{27}\)

OpenStudy (anonymous):

do you know how to do that?

OpenStudy (anonymous):

find a common denominator?

OpenStudy (anonymous):

Basically, what he's saying is that 1/sqrt(a) = sqrt(a) / a. This is because you can multiply by sqrt(a)/sqrt(a) and the root on the bottom cancels out, while the 1 on the top becomes sqrt(a). You get sqrt(12)/12 - sqrt(27)/27, and you know how to simplify from there.

OpenStudy (anonymous):

That's probably a simpler way to look at it, I was going to simplify the denominator and then move it to the top.

OpenStudy (anonymous):

But regardless, you need to simplify the radical.

OpenStudy (anonymous):

but they dont reely have a common denominator do they?

OpenStudy (anonymous):

would it be 3 * 4 and 3 * 9 for it?

OpenStudy (anonymous):

In order to simplify a radical like \(\sqrt{12}\), you need to find a square that divides it. In this case, 4 divides 12, so we write \[\sqrt{12}=\sqrt{4}\sqrt{3}=2\sqrt{3}\]

OpenStudy (anonymous):

Right, exactly.

OpenStudy (anonymous):

Would you prefer to move everything to the top and then deal with it or deal with it first, then move it to the top?

OpenStudy (anonymous):

either way is easiest

OpenStudy (anonymous):

i mean whichever way is easiest

OpenStudy (anonymous):

Let's do it the way quantummodulus suggested. Did you understand what they said?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Ok. So what have you gotten to so far?

OpenStudy (anonymous):

(sqrt 12/2 sqrt 3) - (sqrt 27/3 sqrt 3)

OpenStudy (anonymous):

Ok, not quite. I think we got you jumbled up a bit. Let's take a step back. First, we're going to do what is called rationalizing the denominator. That is where we multiply the top and the bottom by the denominator. This will get rid of the square root. Do you know how to do this?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Ok, so what do you get after rationalizing the denominator?

OpenStudy (anonymous):

i think i square rooted something that didnt need to be sqaure rooted

OpenStudy (anonymous):

so now i have sq rt 12/12 - sqrt 27/27 the denominators not being sqrt

OpenStudy (anonymous):

Yup, perfect. Now, simplify the radical, like you did before.

OpenStudy (anonymous):

on top?

OpenStudy (anonymous):

Yup.

OpenStudy (anonymous):

okay i got 2 sqrt 3/12 - 3 sqrt 3/27

OpenStudy (anonymous):

Right, now cancel and add fractions.

OpenStudy (anonymous):

sqrt 3/6 - sqrt3/9

OpenStudy (anonymous):

Yup, now common denominators and add.

OpenStudy (anonymous):

3/18 - sqrt 6/18

OpenStudy (anonymous):

\(3*\sqrt{3}\) doesn't equal 3, same for the other fraction

OpenStudy (anonymous):

3 sqrt 3/18 - 2sqrt 3/9

OpenStudy (anonymous):

/18

OpenStudy (anonymous):

now just subtract

OpenStudy (anonymous):

sqrt 3/18 is the answer then?

OpenStudy (anonymous):

Yup. So the process is: rationalize the denominator, simplify the radical, simplify fractions

OpenStudy (anonymous):

Sorry about all of the confusion at the beginning.

OpenStudy (anonymous):

oh no that was my fault thank u for all your help thank u!

OpenStudy (anonymous):

You're welcome. Good luck on any other problems.

OpenStudy (anonymous):

\[\sqrt{a}-\sqrt{b} = \sqrt{a-2\sqrt{a*b}+b}\] Plug in 1/12 for a and 1/27 for b and simplify. The result is: \[{\sqrt{1\over 108}} = {1 \over 6\sqrt{3}}\] The formula is obtained by squaring the quantity radical a - radical b, ie: the left side of the formula equation. The first and last terms of the result is a and b respectively. The mid term is negative two times the product of radical a and radical b. Radical a times radical b is equal to the square root of the product of a and b. The final operation is to take the square root of every thing.